OpenStudy (anonymous):

Solve the system x - y = 3 and 3x - 8y = 9. What is the value of y? A. 3 B. 0 C. 6 D. -2

OpenStudy (anonymous):

y=x-3 so 3x-8(x-3)=9 then 3x-24x+24=9 and -21x=-21 so x=1 and y=x-3=-2

OpenStudy (hoblos):

B

OpenStudy (hoblos):

@barbillus 8(x-3)= 8x-24

OpenStudy (anonymous):

oops sorry